OpenStudy (anonymous):

What equation would I use to answer this question A 10 inch TV set also has a width of 8 inches what is its height?

OpenStudy (anonymous):

I get six inches am I correct

OpenStudy (anonymous):

 If you have a triangle with a 90 degree angle, you can see it has a hypotenuse, in your case is the diagonal, and 2 catheti. You know that one of them is 8inch. Now you have to use the Pythagorean Theorem, which says: h^2 = c1^2 + c2^2 where h is hypotenuse and c1, c2 are the catheti. Now plug in the known values: 10^2 = 8^2 + x^2 (call x the unknown cathetus) x^2 = 100 - 64 = 36 x = sqrt(36) = 6 inch